5x+3= -4(1-x)
Solve x

Answer by algebrahouse.com(1659)

5x + 3 = -4(1 - x)
5x + 3 = -4 + 4x {used distributive property}
5x = -7 + 4x {subtracted 3 from both sides}
x = -7 {subtracted 4x from both sides}
www.algebrahouse.com